A marketer of motorbikes determines that in 2017, the demand function for its products is
P=2000 - 50Q
Where P is the price ($) of motorbike, and Q is the number of motorbikes sold per month.
i. To sell 20 motorbikes per month, what price should be charged?
ii. If the motorbikes sell at a price of $500, how many motorbikes will be sold per month?
iii. What is the price elasticity of demand if the price equals $500?
iv. At what price, if any, will the demand for motorbikes be unitary elastic?
1
Expert's answer
2018-08-05T07:21:08-0400
P=2000-50*20=1000
500=2000-50*Q;
Q=30
Q=-0,02P+40
E_d^p=-0,02*500/30=-0,33

-1=-0,02*P/(-0,02P+40)

(-0,02*P)/(-0,02P+40)+(-0,02P+40)/(-0,02P+40)=0
-0,04P+40=0
P=1000